A comparative balance sheet for Monty Corporation is presented as follows. Assets Cash Accounts receivable Inventory Land EquPrepare a statement of cash flows for 2020 for Monty Corporation. (Show amounts that decrease cash flow with either a - sign<> Issued common stock to retire $ of bonds outstandingDetermine Monty Corporations current cash debt coverage, cash debt coverage, and free cash flow. (Round current cash debt co
Answer #1
Monty corporation
Statement of cash flows
For the year ended December 31,2020
$ $
Operating activities:
Net income (Profit for the year) 128560
Adjustments to reconcile net income to
Depreciation expense (Note:1) 26700
Increase in accounts receivable (83780-68080) -15700
Increase in inventory (191080-181780) 9300
Decrease in accounts payable (49080-35780) -13300 7000
Cash flow from operating activities a 135560
Investing activities:
Purchase of equipment (261780-202080) -59700
Sale of land (112080-72780) 39300
Cash flow from investing activities b -20400
Financing activities:
Dividend paid -64460
Cash flow from financing activities c -64460
Net increase in cash a+b+c 50700
Add:Beginning balance of cash 22000
Ending balance of cash 72700
Note:1
Depreciation expense=Ending balance of accumulated depreciation-Beginning balance of accumulated depreciation=70780-44080=$ 26700
Common stock issued=Increase in common stock=214000-164000=$ 50000
Bonds retired=Decrease in bonds payable=200000-150000=$ 50000
Issued common stock to retire $ 50000 of bonds outstanding
Current cash debt coverage ratio=Net cash provided by operating activities/Average current liabilities
Average current liabilities=(Beginning current liabilities+Ending current liabilities)/2
Beginning current liabilities=Accounts payable for 2019=$ 49080
Ending current liabilities=Accounts payable for 2020=$ 35780
Average current liabilities=(49080+35780)/2=$ 42430
Current cash debt coverage ratio=135560/42430=$ 3.19
Cash debt coverage ratio=Net cash provided by operating activities/Average total liabilities
Average total liabilities=(Beginning total liabilities+Ending total liabilities)/2
Beginning total liabilities (2019):
$
Accounts payable 49080
Bonds payable 200000
Total 249080
Ending total liabilities (2020):
$
Accounts payable 35780
Bonds payable 150000
Total 185780
Average total liabilities=(249080+185780)/2=$ 217430
Cash debt coverage ratio=135560/217430=0.62
Free cash flow=Net cash provided by operating activities-Capital expenditures for the year
Capital expenditures for the year=Cash flow used in investing activities+Depreciation expense=20400+26700=$ 47100
Free cash flow=Net cash provided by operating activities-Capital expenditures for the year=135560-47100=$ 88460
Comment:
Monty has excellent liquidity.Its financial flexibility is good